OpenStudy (anonymous):

1.solve for g. √6g - 23 = √12 - g 2.solve for d. √d^2 -11 = 5

OpenStudy (a1234):

I think you need to square each side and then solve it like a regular equation.

OpenStudy (a1234):

That is for the first equation. I don't know about the second one.

OpenStudy (kirbykirby):

Is the equation \(\sqrt{d^2-11}=5\) or \(\sqrt{d^2}-11=5\)

OpenStudy (anonymous):

its the first one

OpenStudy (kirbykirby):

\(\sqrt{d^2-11}=5\) \(\left(\sqrt{d^2-11}\right)^2=5^2\) \(d^2-11=25\) \(d^2=36\) \(d=\pm \sqrt{36}=\pm 6\)

OpenStudy (anonymous):

what about the second one??

OpenStudy (kirbykirby):

\(\sqrt{6g - 23} = \sqrt{12 - g}\) \(\left(\sqrt{6g - 23}\right)^2 = \left(\sqrt{12 - g}\right)^2\) \( 6g-23=12-g\) \(7g=35\) \(g=5\)

OpenStudy (anonymous):

thanks could you help with a couple more

OpenStudy (anonymous):

which is the extraneous solution of -x = √2x + 15

OpenStudy (kirbykirby):

If you mean \(\sqrt{2x+15}\) instead of \(\sqrt{2x}+15\), please at least write √(2x+15) to say that you're taking the square root of all of those terms. Otherwise it is not clear.

OpenStudy (anonymous):

yes its the first one sorry

OpenStudy (kirbykirby):

\[-x=\sqrt{2x+15}\\ (-x)^2=\left(\sqrt{2x+15}\right)^2\\ x^2=2x+15\\ x^2-2x-15=0\\ (x-5)(x-3)=0\\ x=5, \text{or }x=-3\] But now look at the original equation...\(-x=\sqrt{2x+15}\) Does it make sense that x=5? No! because that means \(-5=\sqrt{2x+15}\) and a square root cannot equal a negative number. But it works for x=-3, since you get \(-(-3)=3=\sqrt{2x+15}\) So x=5 is the extraneous solution (meaning it is NOT a valid solution to the original equation)

OpenStudy (kirbykirby):

Oh I made a typo on the before last line.. it should read \((x-5)(x+3)=0\)

OpenStudy (anonymous):

thank you